OpenStudy (anonymous):

is the slope of y-3=3(x+1) answer 3

OpenStudy (zale101):

y-3=3(x+1) Distribute 3(x+1) and add 3 to both sides

OpenStudy (charlotte123):

YES! GREAT JOB! http://www.wolframalpha.com/input/?i=y-3%3D3%28x%2B1%29+slope <---

OpenStudy (charlotte123):

Thank You Zale <3

OpenStudy (the_fizicx99):

Your slope is 3 Where y-y1=M(x-x1) M = 3

OpenStudy (anonymous):

Yes awesome job!!

OpenStudy (zale101):